OpenStudy (anonymous):

find the solution set x^2+5x+4=0

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

x^2+5x+4=0 turns into (x+4)(x+1)=0

OpenStudy (anonymous):

is that the answer?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

(x+4)(x+1)=0 x+4=0 or x+1=0 x=?? or x=??

OpenStudy (anonymous):

-4,0

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

idk then

jimthompson5910 (jim_thompson5910):

if x+1=0 then x = ??

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

so -4,-1

jimthompson5910 (jim_thompson5910):

correct